Can anyone answer this?
There are two spinners, one is divided into thirds with green,
yellow and blue. The second spinner is divided into fourths with
blue, yellow, red and white.
find the theoretical probability of spinning blue on both.

Answer by AnlytcPhil(1806)

Can anyone answer this?
There are two spinners, one is divided into thirds with green,
yellow and blue. The second spinner is divided into fourths with
blue, yellow, red and white.
find the theoretical probability of spinning blue on both.

P(blue on 1st) = 1/3
P(blue on 2nd) = 1/4

Since the events are independent, we can multiply their probabilities

P(blue on 1st AND blue on 2nd) = P(blue on 1st) × P(blue on 2nd) = 

(1/3)×(1/4) = 1/12

Edwin 
AnlytcPhil@aol.com